(Please help a friend of mine with these two problems)
1) 2/3x + 4 = 6
2) C + 1 3/4 = 1 2/3

Answer by stanbon(75887)

1) 2/3x + 4 = 6
(2/3)x = 2
x = 2*(3/2)
x = 3
------------
2) C + 1 3/4 = 1 2/3
C + 7/4 = 5/3
C = (5/3)-(7/4)
C = (20/12) - (21/12)
C = -1/12
==============
Cheers,
Stan H.